OpenStudy (maggy14145):

Reece and Kala are playing a game with a die. As Reece rolls the die, Kala notices that it seems to land on the number 6 a lot. She notes that for 5 out of 10 rolls of the die, the die comes up as a 6. The binomial probability of rolling five 6s out of ten throws is approximately _____. The experimental probability of rolling a 6 is _____. Determine the values that accurately complete the previous two sentences. Then use those values to help choose an answer to the following question. Should Kala doubt whether the die is a fair die?

OpenStudy (maggy14145):

- Kala might doubt whether the die is a fair die, since the experimental probability of rolling that many 6s is much greater than the binomial probability. However, since there were only 10 experiments, the die still may be fair.

OpenStudy (maggy14145):

- Kala should definitely not doubt whether the die is a fair die. The two probabilities are relatively close so the die is most likely to be fair.

OpenStudy (maggy14145):

- Kala should definitely doubt whether the die is a fair since the experimental probability of rolling that many 6s is much greater than the binomial probability of doing so.

OpenStudy (maggy14145):

THOSE ARE THE OPTIONS LISTED ABOVE ^^^

jimthompson5910 (jim_thompson5910):

are you familiar with the binomial probability distribution?

OpenStudy (maggy14145):

not quite.

jimthompson5910 (jim_thompson5910):

this formula doesn't look familiar? \[\Large \left(_n C _k\right)*\left(p\right)^{k}*\left(1-p\right)^{n-k}\]

jimthompson5910 (jim_thompson5910):

If you want, you can use this calculator instead http://stattrek.com/online-calculator/binomial.aspx

OpenStudy (maggy14145):

oh yes I've definitely seen it before but am a little confused on how to apply it to an equation

jimthompson5910 (jim_thompson5910):

So you've dealt with permutations and combinations before?

OpenStudy (maggy14145):

I don't think I've ever heard of those terms ever...but for some reason the formula you sent me looks quite familiar

jimthompson5910 (jim_thompson5910):

So you've seen this formula before? \[\Large _n C _r = \frac{n!}{r!*(n-r)!}\] or no?

OpenStudy (maggy14145):

no never

jimthompson5910 (jim_thompson5910):

ok then it will probably be best to stick with the calculator I posted above

OpenStudy (maggy14145):

ok so how do I apply the calculator to the question I'm asking..?

jimthompson5910 (jim_thompson5910):

there are 6 sides on a die. The probability of landing on any one side is 1/6 = 0.1667 we're doing 10 trials we want x to be equal to 5

jimthompson5910 (jim_thompson5910):

so you'll type 0.1667 into the first box 10 in the second box 5 in the third box then hit "calculate"

jimthompson5910 (jim_thompson5910):

what value pops up next to `Binomial Probability: P(X = 5) ` ?

OpenStudy (maggy14145):

.0130

jimthompson5910 (jim_thompson5910):

so that is the answer to `The binomial probability of rolling five 6s out of ten throws is approximately _____. `

jimthompson5910 (jim_thompson5910):

`The experimental probability of rolling a 6 is _____.` five 6's are rolled out of 10 trials, so 5/10 = 0.5 is the answer for this part

jimthompson5910 (jim_thompson5910):

question: is the die fair? why or why not?

OpenStudy (maggy14145):

am I supposed to answer that question?

jimthompson5910 (jim_thompson5910):

it asks `Should Kala doubt whether the die is a fair die?`

OpenStudy (maggy14145):

would the first option be correct?

jimthompson5910 (jim_thompson5910):

what do you mean by "first option"?

OpenStudy (maggy14145):

Kala might doubt whether the die is a fair die, since the experimental probability of rolling that many 6s is much greater than the binomial probability. However, since there were only 10 experiments, the die still may be fair.

jimthompson5910 (jim_thompson5910):

I'm guessing there are other options? What are they?

OpenStudy (maggy14145):

I listed them on the top. can u not see them? if not I can easily copy/paste them again

jimthompson5910 (jim_thompson5910):

oh right, brainfart ok one sec

jimthompson5910 (jim_thompson5910):

I would go with `- Kala should definitely doubt whether the die is a fair since the experimental probability of rolling that many 6s is much greater than the binomial probability of doing so.` because the die seems to be unfair. It seems like there is more than one face with '6' on it. Or the face opposite '6' is weighed down to make 6 come up more frequently

OpenStudy (maggy14145):

ok I was gonna say it was either gonna be she might doubt or just all together doubt the fact that the die isn't fair

OpenStudy (maggy14145):

Thank you again for your help!

jimthompson5910 (jim_thompson5910):

That's a good point. 10 trials may not be enough. Still, I think it is rigged in favor of '6'

OpenStudy (maggy14145):

Also on this question I just wanted to make sure I got the right answer. I'm pretty sure it's the correct one so if you could please take a second to check, that would be amazing

OpenStudy (maggy14145):

Which of the equations of hyperbolas below has a vertical asymptote at x = -3?

OpenStudy (maggy14145):

y=6/(x-4) -3

OpenStudy (maggy14145):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you only listed one equation

OpenStudy (maggy14145):

oh it was either that one or this one

OpenStudy (maggy14145):

y=4/(x-3) +6

jimthompson5910 (jim_thompson5910):

so only two options?

OpenStudy (maggy14145):

yes

jimthompson5910 (jim_thompson5910):

none of the answer choices are the answer. You need a `x+3` in the denominator to have a vertical asymptote of x = -3

OpenStudy (maggy14145):

oh that was my bad! the plus was supposed to be a minus!

jimthompson5910 (jim_thompson5910):

so instead of `y=4/(x-3) +6` it's actually `y=4/(x+3) +6` ?